Solve the following linear programming problems using simplex method.
Maximize Z = 3x+4y subject x+y=4, 2x+y = 5,x=0,y=0.

Answer by ikleyn(52781)

.
Solve the following linear programming problems using simplex method.
Maximize Z = 3x+4y subject x+y=4, 2x+y = 5,x=0,y=0.
~~~~~~~~~~~~~~~~~


            Hello, your post  INCORRECTLY  presents the problem.

            The correct formulation should use  INEQUALITIES  everythere where you write equations.


The correct formulation is THIS


    Solve the following linear programming problems using simplex method.
 
    Maximize Z = 3x+4y 

    subject x+y <= 4, 2x+y <= 5, x >= 0,y >= 0.


Since you presented the problem incorrectly, I will not work for you
until you will learn from a right source what you want and what you need.